### v2

@@ 2.4 SCD mappings
${\mathcal{S}}F(x,y):$ =$=\{L\in{\cal Z}_{n,m}\,\big{|}\,\exists(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}(x,y)\mbox{ such that }\lim_{k\to\infty}d_{{\cal Z}_{n,m}}\big{(}L,T_{\mathrm{gph}\,F}(x_{k},y_{k})\big{)}=0\},$
@@
By virtue of the definition, the SCD property at $(\bar{x},\bar{y})$ is obviously equivalent with the condition ${\mathcal{S}}F(\bar{x},\bar{y})\not=\emptyset$. Since we consider convergence in the compact metric space ${\cal Z}_{n,m}$ we readily obtain that $F$ has the SCD property at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ if and only if $(\bar{x},\bar{y})\in\mathrm{cl}\,{\cal O}_{F}$.
@@
The derivatives ${\mathcal{S}}F$ and ${\mathcal{S}}^{*}F$ can be considered as a generalization of the B-Jacobian to multifunctions. In case of single-valued continuous mappings one has the following relationship.
Lemma 2.9 (cf. [13, Lemma 3.5]).
Let $\Omega\subset\mathbb{R}^{n}$ be open and let $f:\Omega\to\mathbb{R}^{m}$ be continuous. Then for every $x\in\Omega$ there holds
@@
If $f$ is Lipschitz near $x$, these inclusions hold with equality and $f$ has the SCD property at all points $x^{\prime}$ sufficiently close to $x$.
@@
${\mathcal{S}}F(\bar{x},\bar{y}$ )=$)=\{\nabla\Phi(\bar{x},\bar{y})^{-1}{\rm rge\;}(I,A)\,\big{|}\,A\in\overline{\nabla}f(\bar{u})\},$
${\mathcal{S}}^{*}F(\bar{x},\bar{y})$ =$=\{S_{nm}\nabla\Phi(\bar{x},\bar{y})^{T}S^{T}_{nm}{\rm rge\;}(I,A^{T})\,\big{|}\,A\in\overline{\nabla}f(\bar{u})\}\mbox{ with }S_{nm}:=\begin{pmatrix}0&-I_{m}\\ I_{n}&0\end{pmatrix},$
@@
Prominent examples for graphically Lipschitzian mappings are the subdifferential mapping $\partial f$ for prox-regular and subdifferentially continuous functions $f:\mathbb{R}^{n}\to(-\infty,\infty]$, cf. [35, Proposition 13.46] or, more generally, locally maximally hypomonotone mappings, cf. [12].
@@ 2.5 On semismooth mappings
A locally Lipschitz mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, is semismooth at $\bar{x}$ if and only if $F$ is G-semismooth at $\bar{x}$ [11, Proposition 3.7]. However, there are also single-valued mappings which are semismooth but neither locally Lipschitz and nor G-semismooth, e.g., the function $x\mapsto\sqrt{|x|}$ at 00.
@@
We say that $F$ is $\Gamma$-ss around $(\bar{x},\bar{y})$ if $F$ is $\Gamma$-ss for all $(x,y)\in\mathrm{gph}\,F$ sufficiently close to $(\bar{x},\bar{y})$. Finally, $F$ is called $\Gamma$-ss if $F$ is $\Gamma$-ss at every $(x,y)\in\mathrm{gph}\,F$.
@@
We say that the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is SCD-ss at (around) $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ if it is semismooth at (around) $(\bar{x},\bar{y})$ with respect to the mapping $\bigcup{\mathcal{S}}^{*}F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}\times\mathbb{R}^{n}$ given by
@@
The mapping $F$ is called SCD-ss if it is SCD-ss at every point of its graph.
It follows from the definition of the $\Gamma$-ss property that $F$ has the SCD property around $(\bar{x},\bar{y})$ whenever $F$ is SCD-ss at $(\bar{x},\bar{y})$.
The SCD-ss property has also a useful primal interpretation:
@@ 3 On $\Psi$-semismooth single-valued mappings
The following definition is closely related to the definition of semismooth operators in Banach spaces, cf. [38, Definition 3.1].
@@
    $\lim_{{\textstyle{x\mathop{\rightarrow}\limits^{\Omega}\bar{x}\atop x\not=\bar{x}}}}\frac{\sup\{\|F(x)-F(\bar{x})-A(x-\bar{x})\|\,\big{|}\,A\in\Psi(x)\}}{\|x-\bar{x}\|}=0.$
@@
Our definition of semismoothness differs from Ulbrich’s one [38, Definition 3.1] insofar as we do not suppose local continuity of $F$ but require local boundedness of $\Psi$. If $F$ is $\Psi$-ss at $\bar{x}$ then $\Psi$ is also called a Newton map for $F$ at $\bar{x}$ in [20]. If $\Psi$ is single-valued then it is also called a Newton derivative (or slanting function) for $F$ at $\bar{x}$ in the literature. We dispense with using this terminology because they are normally used at single points, namely solutions of the nonsmooth equation $F(x)=0$. However, for our purposes it is essential that the properties hold not only at one single point $\bar{x}$ but also on a neighborhood of $\bar{x}$.
@@
This construction resembles the one of Qi’s C-subdifferential, see, e.g. [38, Proposition 2.1].
Continuously differentiable mappings $F$ are $\nabla F$-ss. Further, it is easy to see that a mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, which is $\Psi$-ss at $\bar{x}\in\Omega$, is calm at $\bar{x}$ and $\mathrm{clm}\,F(\bar{x})\leq\mathrm{bnd}\,\Psi(\bar{x})$.
@@
Consider an open set $\Omega\subset\mathbb{R}^{n}$, two mappings $F_{i}:\Omega\to\mathbb{R}^{m}$ and two mappings $\Psi_{i}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$, $i=1,2$. If $F_{i}$ is $\Psi_{i}$-ss, $i=1,2$, at $\bar{x}\in\Omega$ then $F_{1}+F_{2}$ is $(\Psi_{1}+\Psi_{2})$-ss at $\bar{x}$.
@@
    For every $x\in\Omega$ there holds $\mathrm{conv}\,\overline{\nabla}F(x)\subset({\rm cocl}\,\Psi)(x)$.
@@
ad (ii): Let $\Pi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$ denote the mapping which assigns to each $x$ the set of all $m\times n$ matrices whose $i$-th row belongs to $({\rm cocl}\,\Psi_{i})(x)$. Clearly, $({\rm cocl}\,\Psi)(x)\subset\Pi(x)$, $x\in\mathbb{R}^{n}$. Next let $\tilde{\Omega}:=\bigcap_{i=i}^{m}\Omega_{\Psi_{i}}$. For every $x\in\tilde{\Omega}$ we obtain from Proposition 3.7(ii) that for every $i=1,\ldots,m$ the component function $F_{i}$ is strictly differentiable at $x$ and $({\rm cocl}\,\Psi_{i})(x)=\{\nabla F_{i}(x)\}$ is a singleton. Hence, $F$ is also strictly differentiable at $x$ and $({\rm cocl}\,\Psi)(x)=\Pi(x)=\{\nabla F(x)\}$. Since the set $\Omega\setminus\tilde{\Omega}=\bigcup_{i=1}^{m}(\Omega\setminus\Omega_{\Psi_{i}})$ is negligible, the set $\tilde{\Omega}$ differs from the set ${\cal O}_{F}$, the set of all points where $F$ is differentiable, by a negligible set. By defining the mapping $\nabla_{s}F(x):=\nabla F(x)$ for $x\in\mathrm{dom}\,\nabla_{s}F:=\tilde{\Omega}$ and taking into account $\mathrm{conv}\,\overline{\nabla}F={\rm cocl}\,\nabla F$, it follows from [39, Theorem 4] that $\mathrm{conv}\,\overline{\nabla}F={\rm cocl}\,\nabla_{s}F$. Since $\nabla_{s}F(x)\subset({\rm cocl}\,\Psi)(x)$ $\forall x\in\Omega$, we obtain $\mathrm{conv}\,\overline{\nabla}F\subset\big{(}{\rm cocl}\,({\rm cocl}\,\Psi)\big{)}(x)=({\rm cocl}\,\Psi)(x)$ $\forall x\in\Omega$.
@@
4 On strict proto-differentiability and the SCD-ss property of graphically Lipschitzian mappings
@@
$\bar{L}\supset\mathop{{\rm Lim}\,{\rm sup}}_{(x,y){\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)\supset\bigcup_{L\in{\mathcal{S}}F(\bar{x},\bar{y})}L\not=\emptyset.$
@@
11 11 $=\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}w_{i}\|\leq\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}w_{i}^{k}\|+\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}(w_{i}-w_{i}^{k})\|$
@@
Thus, $F$ is strictly proto-differentiable at $(\bar{x},\bar{y})$ if and only if $f$ is strictly proto-differentiable at $(\bar{u},f(\bar{u}))$, which is the same as $f$ being strictly differentiable at $\bar{u}$. Now, by Lemma 2.11 the SC derivative ${\mathcal{S}}F(\bar{x},\bar{y})$ is a singleton if and only if ¯$\overline{\nabla}f(\bar{u})$ is a singleton, which is equivalent to strict differentiability of $f$ at $\bar{u}$. Combining these statements it follows that $F$ is strictly proto-differentiable at $(\bar{x},\bar{y})$ if and only if ${\mathcal{S}}F(\bar{x},\bar{y})$ is a singleton. ∎
For graphically Lipschitzian mappings, the SCD-ss property is closely related with the G-semismoothness.
@@
Assume that the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ and let $\Phi$ and $f$ be as in Definition 2.10. Then $F$ is SCD-ss at $(\bar{x},\bar{y})$ if and only if $f$ is ¯$\overline{\nabla}f$-semismooth at $\bar{u}$, where $(\bar{u},f(\bar{u}))=\Phi(\bar{x},\bar{y})$.
@@
and the assertion follows from Proposition 2.15 and the definition of ¯$\overline{\nabla}f$-semismoothness. ∎
@@
Assume that the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ and SCD-ss around $(\bar{x},\bar{y})$. Then there is an open neighborhood $W$ of $(\bar{x},\bar{y})$ such that for almost all $(x,y)\in\mathrm{gph}\,F\cap W$ (with respect to the $n$-dimensional Hausdorff measure) the mapping $F$ is strictly proto-differentiable at $(x,y)$ and ${\mathcal{S}}F(x,y)$ is a singleton.
@@
According to the notation of Definition 2.10, let $\Omega_{F}$ denote the set of all points in $\mathrm{gph}\,F\cap W$ where $F$ is strictly proto-differentiable and let $\Omega_{f}$ denote the set of all points $u\in U\subset\mathbb{R}^{n}$ where $f$ is strictly differentiable. We may also assume that $F$ is SCD-ss on $\mathrm{gph}\,F\cap W$ and therefore $f$ is ¯$\overline{\nabla}f$-ss on $U$ by Proposition 4.2. We conclude from Theorem 3.8 that the set $U\setminus\Omega_{f}$ has Lebesgue measure zero and the assertion follows from the relation
@@
which holds true by Theorem 4.1 and Lemmas 2.1, 2.11. ∎

### v3

@@ 2.4 SCD mappings
${\mathcal{S}}F(x,y):$ $=\{L\in{\cal Z}_{n,m}\,\big|\,\exists(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}(x,y)\mbox{ such that }\lim_{k\to\infty}d_{{\cal Z}_{n,m}}\big(L,T_{\mathrm{gph}\,F}(x_{k},y_{k})\big)=0\},$
@@
By virtue of the definition, the SCD property at $(\bar{x},\bar{y})$ is obviously equivalent with the condition ${\mathcal{S}}F(\bar{x},\bar{y})\not$$\emptyset$. Since we consider convergence in the compact metric space ${\cal Z}_{n,m}$ we readily obtain that $F$ has the SCD property at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ if and only if $(\bar{x},\bar{y})\in\mathrm{cl}\,{\cal O}_{F}$.
@@
The derivatives ${\mathcal{S}}F$ and ${\mathcal{S}}^{*}F$ can be considered as a generalization of the B-Jacobian to multifunctions. If $f:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, is continuous, then for every $x\in\Omega$ there holds
@@
If $f$ is Lipschitz near $x$, these inclusions hold with equality and $f$ has the SCD property at all points $x^{\prime}$ sufficiently close to $x$, cf.[13, Lemma 3.5].
@@
${\mathcal{S}}F(\bar{x},\bar{y}$ $)=\{\nabla\Phi(\bar{x},\bar{y})^{-1}{\rm rge\;}(I,A)\,\big|\,A\in\overline{\nabla}f(\bar{u})\},$
${\mathcal{S}}^{*}F(\bar{x},\bar{y})$ $=\{S_{nm}\nabla\Phi(\bar{x},\bar{y})^{T}S^{T}_{nm}{\rm rge\;}(I,A^{T})\,\big|\,A\in\overline{\nabla}f(\bar{u})\}\mbox{ with }S_{nm}:=\begin{pmatrix}0&-I_{m}\\ I_{n}&0\end{pmatrix}$,
@@
Prominent examples for graphically Lipschitzian mappings are the subdifferential mapping $\partial f$ for prox-regular and subdifferentially continuous functions $f:\mathbb{R}^{n}\to(-\infty,\infty]$, cf. [36, Proposition 13.46] or, more generally, locally maximally hypomonotone mappings, cf. [12].
@@ 2.5 On semismooth mappings
A locally Lipschitz mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, is semismooth at $\bar{x}$ if and only if $F$ is G-semismooth at $\bar{x}$ [11, Proposition 3.7]. However, there are also single-valued mappings which are semismooth but neither locally Lipschitz and nor G-semismooth, e.g., the function $x\mapsto\sqrt{|x|}$ at 0.
@@
We say that $F$ is $\Gamma$-$s$ around $(\bar{x},\bar{y})$ if $F$ is $\Gamma$-$s$ for all $(x,y)\in\mathrm{gph}\,F$ sufficiently close to $(\bar{x},\bar{y})$. Finally, $F$ is called $\Gamma$-ss if $F$ is $\Gamma$-ss at every $(x,y)\in\mathrm{gph}\,F$.
@@
We say that the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is SCD-$s$ at (around) $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ if it is semismooth at (around) $(\bar{x},\bar{y})$ with respect to the mapping $\bigcup{\mathcal{S}}^{*}F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}\times\mathbb{R}^{n}$ given by
@@
The mapping $F$ is called SCD-$s$ if it is SCD-$s$ at every point of its graph.
It follows from the definition of the $\Gamma$-ss property that $F$ has the SCD property around $(\bar{x},\bar{y})$ whenever $F$ is SCD-$s$ at $(\bar{x},\bar{y})$.
The SCD-$s$ property has also a useful primal interpretation:
@@ 3 On $\Psi$-semismooth single-valued mappings
The following definition is closely related to the definition of semismooth operators in Banach spaces, cf. [39, Definition 3.1].
@@
    $\lim_{{\textstyle{x\mathop{\rightarrow}\limits^{\Omega}\bar{x}\atop x\not=\bar{x}}}}\frac{\sup\{\|F(x)-F(\bar{x})-A(x-\bar{x})\|\,\big|\,A\in\Psi(x)\}}{\|x-\bar{x}\|}=0.$ (3.10)
@@
Our definition of semismoothness differs from Ulbrich’s one [39, Definition 3.1] insofar as we do not suppose local continuity of $F$ but require local boundedness of $\Psi$. If $F$ is $\Psi$-ss at $\bar{x}$ then $\Psi$ is also called a Newton map for $F$ at $\bar{x}$ in [20]. If $\Psi$ is single-valued then it is also called a Newton derivative (or slanting function) for $F$ at $\bar{x}$ in the literature. We dispense with using this terminology because they are normally used at single points, namely solutions of the nonsmooth equation $F(x)=0$. However, for our purposes it is essential that the properties hold not only at one single point $\bar{x}$ but also on a neighborhood of $\bar{x}$. For instance, for solving (1.1) with the BT-algorithm, we will show below that any semismooth derivative of $\theta$ on an open superset of $U_{\rm ad}$ is a feasible choice for the oracle.
@@
Though we do not explicitly require (Lipschitz) continuity of $F$ in Definition 3.1 , it is implied by (3.10) together with the boundedness of $\Psi$. It is easy to see that a mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, which is $\Psi$-ss at $\bar{x}\in\Omega$, is calm at $\bar{x}$ and $\mathrm{clm}\,F(\bar{x})\leq\mathrm{bnd}\,\Psi(\bar{x})$. Further, as we will show below in Theorem 3.8, if $F$ is $\Psi$-ss on $\Omega$, then it is also locally Lipschitz there. The main reason for omitting the requirement of Lipschitz continuity in Definition 3.1 is that we want to deal also with mappings which are implicitly defined by solution mappings of generalized equations, where verifying Lipschitz continuity with conventional tools like limiting coderivatives might be a difficult issue.
There is a big difference between pointbased semismoothnes and semismoothness on a neighborhood. For instance, given any mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, and $\bar{x}\in\Omega$, the mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$ given by $\mathrm{dom}\,\Psi=\Omega$ and
$\Psi(x)=\begin{cases}\frac{(F(x)-F(\bar{x}))(x-\bar{x})^{T}}{\|x-\bar{x}\|^{2}}&\mbox{if $x\in\Omega\setminus\{\bar{x}\}$,}\\ 0&\mbox{if $x=\bar{x}$}\end{cases}$
obviously satisfies (3.10). If $F$ is calm at $\bar{x}$ then $\Psi$ is also locally bounded at $\bar{x}$ and therefore $F$ is $\Psi$-ss at $\bar{x}$. However, $F$ will not be $\Psi$-ss on any neighborhood of $\bar{x}$, except for very special cases. As an example consider $F:\mathbb{R}\to\mathbb{R}$, $F(x):=x^{2}$ and $\bar{x}:=0$. The formula above gives $\Psi(x)=x$ and
$\lim_{{\textstyle{x\to x^{\prime}\atop x\not=x^{\prime}}}}\frac{\|F(x)-F(x^{\prime})-\Psi(x)(x-x)\|}{\|x-x^{\prime}\|}=\lim_{{\textstyle{x\to x^{\prime}\atop x\not=x^{\prime}}}}\frac{|x^{2}-x^{\prime 2}-x(x-x^{\prime})|}{|x-x^{\prime}|}=|x^{\prime}|\not=0$
for every $x^{\prime}\not=\bar{x}$.
Moreover, in the pointbased setting we may modify a semismooth derivative $\Psi$ by adding matrices of order ${\cal O}(\|F(x)-F(\bar{x})\|)$ to $\Psi(x)$ without violating (3.10). This issue is sometimes exploited in semismooth Newton methods for finding zeros of $F$ by neglecting terms of order ${\cal O}(\|F(x)\|)$, see, e.g. [32]. However, with such a modification the semismoothness property on a neighborhood will no longer hold.
@@
This construction resembles the one of Qi’s C-subdifferential, see, e.g. [39, Proposition 2.1].
Clearly, continuously differentiable mappings $F$ are $\nabla F$-$s$.
@@
Consider an open set $\Omega\subset\mathbb{R}^{n}$, two mappings $F_{i}:\Omega\to\mathbb{R}^{m}$ and two mappings $\Psi_{i}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$, $i=1,2$. If $F_{i}$ is $\Psi_{i}$-$s$, $i=1,2$, at $\bar{x}\in\Omega$ then $F_{1}+F_{2}$ is $(\Psi_{1}+\Psi_{2})$-ss at $\bar{x}$.
@@
    For every $x\in\Omega$ there holds $\emptyset\not=\mathrm{conv}\,\overline{\nabla}F(x)\subset({\rm cocl}\,\Psi)(x)$.
@@
ad (ii): Let $\Pi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$ denote the mapping which assigns to each $x$ the set of all $m\times n$ matrices whose $i$-th row belongs to $({\rm cocl}\,\Psi_{i})(x)$. Clearly, $({\rm cocl}\,\Psi)(x)\subset\Pi(x)$, $x\in\mathbb{R}^{n}$. Next let $\tilde{\Omega}:=\bigcap_{i=i}^{m}\Omega_{\Psi_{i}}$. For every $x\in\tilde{\Omega}$ we obtain from Proposition 3.7(ii) that for every $i=1,\ldots,m$ the component function $F_{i}$ is strictly differentiable at $x$ and $({\rm cocl}\,\Psi_{i})(x)=\{\nabla F_{i}(x)\}$ is a singleton. Hence, $F$ is also strictly differentiable at $x$ and $({\rm cocl}\,\Psi)(x)=\Pi(x)=\{\nabla F(x)\}$. Since the set $\Omega\setminus\tilde{\Omega}=\bigcup_{i=1}^{m}(\Omega\setminus\Omega_{\Psi_{i}})$ is negligible, the set $\tilde{\Omega}$ differs from the set ${\cal O}_{F}$, the set of all points where $F$ is differentiable, by a negligible set. By defining the mapping $\nabla_{s}F(x):=\nabla F(x)$ for $x\in\mathrm{dom}\,\nabla_{s}F:=\tilde{\Omega}$ and taking into account $\mathrm{conv}\,\overline{\nabla}F={\rm cocl}\,\nabla F$, it follows from [40, Theorem 4] that $\mathrm{conv}\,\overline{\nabla}F={\rm cocl}\,\nabla_{s}F$. Since $\nabla_{s}F(x)\subset({\rm cocl}\,\Psi)(x)$ $\forall x\in\Omega$, we obtain $\mathrm{conv}\,\overline{\nabla}F\subset\big({\rm cocl}\,({\rm cocl}\,\Psi)\big)(x)=({\rm cocl}\,\Psi)(x)$ $\forall x\in\Omega$.
@@
4 On strict proto-differentiability and the SCD-$s$ property of graphically Lipschitzian mappings
@@
$\bar{L}\supset\mathop{{\rm Lim}\,{\rm sup}}_{(x,y){\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)\supset\bigcup_{L\in{\mathcal{S}}F(\bar{x},\bar{y})}L\not$$\emptyset$.
@@
$1$ $=\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}w_{i}\|\leq\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}w_{i}^{k}\|+\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}(w_{i}-w_{i}^{k})\|$
@@
Thus, $F$ is strictly proto-differentiable at $(\bar{x},\bar{y})$ if and only if $f$ is strictly proto-differentiable at $(\bar{u},f(\bar{u}))$, which is the same as $f$ being strictly differentiable at $\bar{u}$. Now, by Lemma 2.8 the SC derivative ${\mathcal{S}}F(\bar{x},\bar{y})$ is a singleton if and only if $\overline{\nabla}f(\bar{u})$ is a singleton, which is equivalent to strict differentiability of $f$ at $\bar{u}$. Combining these statements it follows that $F$ is strictly proto-differentiable at $(\bar{x},\bar{y})$ if and only if ${\mathcal{S}}F(\bar{x},\bar{y})$ is a singleton. ∎
For graphically Lipschitzian mappings, the SCD-$s$ property is closely related with the G-semismoothness.
@@
Assume that the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ and let $\Phi$ and $f$ be as in Definition 2.7. Then $F$ is SCD-$s$ at $(\bar{x},\bar{y})$ if and only if $f$ is $\overline{\nabla}f$-semismooth at $\bar{u}$, where $(\bar{u},f(\bar{u}))=\Phi(\bar{x},\bar{y})$.
@@
and the assertion follows from Proposition 2.12 and the definition of $\overline{\nabla}f$-semismoothness. ∎
@@
Assume that the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ and SCD-$s$ around $(\bar{x},\bar{y})$. Then there is an open neighborhood $W$ of $(\bar{x},\bar{y})$ such that for almost all $(x,y)\in\mathrm{gph}\,F\cap W$ (with respect to the $n$-dimensional Hausdorff measure) the mapping $F$ is strictly proto-differentiable at $(x,y)$ and ${\mathcal{S}}F(x,y)$ is a singleton.
@@
According to the notation of Definition 2.7, let $\Omega_{F}$ denote the set of all points in $\mathrm{gph}\,F\cap W$ where $F$ is strictly proto-differentiable and let $\Omega_{f}$ denote the set of all points $u\in U\subset\mathbb{R}^{n}$ where $f$ is strictly differentiable. We may also assume that $F$ is SCD-$s$ on $\mathrm{gph}\,F\cap W$ and therefore $f$ is $\overline{\nabla}f$-ss on $U$ by Proposition 4.2. We conclude from Theorem 3.8 that the set $U\setminus\Omega_{f}$ has Lebesgue measure zero and the assertion follows from the relation
@@
which holds true by Theorem 4.1 and Lemmas 2.1, 2.8. ∎